If z = 2 ( cos 2Â¥/3 + isin 2Â¥/3 ), then z^3=


i. 2 - i8√3

ii. -4 + i4√3

iii.-8-i√2

iv. -8 - i4


1
Expert's answer
2021-08-24T08:38:21-0400

If "Z=2(cos(\frac{2¥}{3})+isin(\frac{2¥}{3}))"

we know de moivre's formula

:-

"Z^n=r(re^{r\theta})^n=r^n(cosn\theta+isinn\theta)\\orZ^n=r^n[cos(n¥)+isin(n¥)]\\so,\space Z^3=2^3(cos(\frac{2¥}{3})3+isin(\frac{2¥}{3}.3))\\=8(cos(2¥)+isin(2¥))\\=8(\frac{-1}{2}+i.\frac{\sqrt3}{2})"